Bond prices depend on the market rate of interest, stated rate of interest, and time. Read the requirements Requirement 1. Compute the price of the following 7% bonds of Friendship Telecom. a. The price of the $200,000 bond issued at 74.50 is $ The price of the $200,000 bond issued at 104.50 is s b. T c. The price of the $200,000 bond issued at 94.75 is $ c. The price of the $200,000 bond issued at 102.25 is $ Requirement 2. Which bond will Friendship Telecom have to pay the most to retire at maturity? Explain your answer Bond a. because it was issued at the lowest price Bond b. because it was issued at the highest price. Bond c. because it was issued at a discount. Bond d. because it was issued at a premium. Friendship Telecom will pay $200,000 at maturity for all four of the bonds. The bonds all have the same maturity value
Answer #1

Requirement 1

a. The price of the 200000 bond issued at 74.50 is

$ 1,49,000.00

a. The price of the 200000 bond issued at 104.50 is

$ 2,09,000.00

a. The price of the 200000 bond issued at 94.75 is

$ 1,89,500.00

a. The price of the 200000 bond issued at 102.25 is

$ 2,04,500.00

Working

a. The price of the 200000 bond issued at 74.50 is $

=200000/100*74.5

a. The price of the 200000 bond issued at 104.50 is $

=200000/100*104.5

a. The price of the 200000 bond issued at 94.75 is $

=200000/100*94.75

a. The price of the 200000 bond issued at 102.25 is $

=200000/100*102.25

Requirement 2

Correct answer—Friendship telecom will pay $200000 at maturity for all four of the bonds. The bonds all have the same maturity value.

Bonds are issued at premium, discount or at par but redemption is done at Face value.